gelphielvr:

Does anyone know how I can calculate the mole faster? Like barium chloride --> BaCl --> their respective numbers ---> the result g/mol like is there a faster way to calculate it or is this the only way?

ihy:

Write the formula – BaCl₂ Count each element – Ba = 1, Cl = 2 Multiply by atomic mass (from periodic table) Ba ≈ 137.33 g/mol Cl ≈ 35.45 g/mol → ×2 = 70.90 g/mol Add → 137.33 + 70.90 = 208.23 g/mol

gelphielvr:

@ihy wrote:
Write the formula – BaCl₂ Count each element – Ba = 1, Cl = 2 Multiply by atomic mass (from periodic table) Ba ≈ 137.33 g/mol Cl ≈ 35.45 g/mol → ×2 = 70.90 g/mol Add → 137.33 + 70.90 = 208.23 g/mol
ok thanks but do you know if there's a way to do it faster?

gelphielvr:

@ihy wrote:
Write the formula – BaCl₂ Count each element – Ba = 1, Cl = 2 Multiply by atomic mass (from periodic table) Ba ≈ 137.33 g/mol Cl ≈ 35.45 g/mol → ×2 = 70.90 g/mol Add → 137.33 + 70.90 = 208.23 g/mol
wait why did you multiply it?

gelphielvr:

ohh I got it

xXAikoXx:

start with the formula BaCl₂ and take it from there (:

gelphielvr:

@xxaikoxx wrote:
start with the formula BaCl₂ and take it from there (:
very helpful .